# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

### v1

@@ 3.2.1 Autoregressive Transformer Architecture
We initialize $g_{i}=0$ to retain VAR’s pretrained behavior and gradually introduce conditioning. Furthermore, we replace absolute positional embeddings in VAR with 2D Rotary Positional Embeddings (RoPE) for scaling resolution from $256\times 256$ to $512\times 512$, as RoPE is well-suited for handling varying sequence lengths rope . We also remove AdaLN layers, reducing $\sim 100$M parameters with negligible impact on performance. Inference closely follows that of VAR (see Sec. 3.1), except that each scale prediction is now guided by the degraded latent. The output is a sequence of predicted indices $\{\hat{r}^{k}_{\text{pred}}\}_{k=1}^{K}$, which is then used to construct the discrete restored latent $f^{\text{pred}}_{\text{quant}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. The above steps are shown in Fig. 2(a)(ii). More architectural details are given in the supplementary.
@@ 3.2.2 Detail-Preserving Restoration
The discrete latent ($f^{\text{pred}}_{\text{quant}}$) predicted by the RestoreVAR transformer is decoded by the VQVAE to produce the restored image. However, vector-quantization and VAE decoding cause a noticeable loss of fine details in the pixel-space, leading to distorted reconstructions (see Fig. 4(b)). This presents a major challenge for using VAR in AiOR, as the scene semantics may not be accurately preserved. To address this, we introduce VAE decoder fine-tuning on continuous latents, and a lightweight latent refinement transformer (LRT) that converts discrete latents to continuous latents for decoding.
VAE Decoder Fine-Tuning. HART hart addressed VAE-induced distortions by fine-tuning the VAE decoder on both discrete and continuous latents. While effective for generative tasks, the VAE decoder of HART produces overly textured outputs, compromising accurate reconstruction (see Fig. 4(c)). Instead, we fine-tune the decoder only on continuous latents, bypassing the quantizer. The encoder and quantizer are kept frozen, and the decoder is trained on $(f^{\text{gt}}_{\text{cont}},I_{\text{gt}})$ pairs. To avoid overly smooth outputs, we use a PatchGAN patchgan discriminator (see Sec. 4.4) and optimize the decoder using pixel-wise, perceptual, and adversarial losses as
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},$
(a) Input
(b) VAR
(c) HART
(d) Ours
Figure 3: Qualitative comparisons of the input reconstructed using VAR var , HART hart and Our VAE decoder. Our result has minimal distortions.
@@
where $\mathcal{L}_{\text{L1}}$ is the L1 loss, $\mathcal{L}_{\text{SSIM}}$ is the SSIM loss, $\mathcal{L}_{\text{percep}}$ is the perceptual loss, $\mathcal{L}_{\text{adv}}$ is the adversarial loss and $\lambda_{i}$ are their respective weights (see Fig. 2(a)(iv)). Our fine-tuning approach yields a decoder that is well-aligned with the objectives of AiOR, achieving mean (over $1000$ samples) reconstruction PSNR/SSIM scores of $28.14$dB/$0.842$, outperforming both the VAR VQVAE ($22.59$dB/$0.679$) and HART decoders ($26.48$dB/$0.804$). Fig. 4(d) shows that the reconstruction using our decoder is significantly better than VAR and HART (see zoomed patch).
Refining Discrete Latents. Since the VAE decoder is fine-tuned for continuous latents, the predicted discrete latent, $f^{\text{pred}}_{\text{quant}}$, must be converted into a continuous form for decoding. While HART uses a $37$M parameter diffusion-based MLP for this, it incurs a $\sim 20\%$ inference overhead due to iterative denoising. Instead, we propose a lightweight, non-generative latent refinement transformer (LRT) that predicts a residual, which when added to $f^{\text{pred}}_{\text{quant}}$, produces a continuous latent, $\hat{f}_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$ as
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),$
@@ 4.1 Implementation Details
Each component of RestoreVAR was trained independently to disentangle learning objectives. We used the VAR model of depth $16$ as the RestoreVAR transformer backbone and trained it with the AdamW optimizer adamw , a learning rate (LR) of $10^{-4}$, batch size of $48$, for $100$ epochs. The latent refiner was trained for $100$ epochs with the AdamW optimizer, LR$=10^{-4}$ and a batch size of $96$. The VAE decoder was fine-tuned using a weighted combination of losses (see Sec. 3.2.2) with empirically chosen weights: $\lambda_{1}=2.0$, $\lambda_{2}=0.4$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.01$. Fine-tuning was performed for $5$ epochs with a learning rate of $3\times 10^{-4}$ and a batch size of $12$, using AdamW. All training was conducted on $8$ RTX A6000 GPUs, while inference was performed on a single RTX 4090 GPU.
@@ 4.2 Datasets
We trained RestoreVAR for five restoration tasks: dehazing, desnowing, deraining, low-light enhancement and deblurring. For dehazing, we used the RESIDE reside dataset comprising $72135$ training and $500$ test images. The Snow100k dataset snow100k was used for desnowing, with 50,000 training images and 16,801 test images (heavy subset). For deraining, we used Rain13K mprnet consisting of 13,711 training images and 4,298 test images. The LOLv1 lolv1 dataset was used for low-light enhancement, consisting of $485$ training images and $15$ test images. For deblurring, we used the GoPro gopro dataset comprising 2,103 training images and 1,111 test images. We additionally test the generalization capability of our model on real-world datasets, namely, LHP lhprain ($1000$ images), REVIDE revide ($284$ images), SICE sice ($565$ images), TOLED toledpoled ($30$ images) and POLED toledpoled ($30$ images).
4.3 Comparisons
We compare RestoreVAR with recent state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR promptir , InstructIR instructir , AWRaCLe awracle and DCPT dcpt . Among generative methods, we compare with the latent diffusion model (LDM) based approaches Diff-Plugin diffplugin , AutoDIR autodir and PixWizard pixwizard . To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for all our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. We were unable to re-train InstructIR and DCPT as they have not released their training codes. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 1 presents quantitative results (PSNR, SSIM and LPIPS scores) on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR significantly outperforms LDM-based AiOR methods across all metrics. These improvements are achieved at a fraction of the computational cost of LDM-based approaches (see Table 5), highlighting the efficacy of our framework. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works diffplugin ; autodir ; awracle ), and assess perceptual realism through a user study.
@@ 4.3 Comparisons
Method Venue RESIDE reside Snow100k snow100k Rain13K mprnet LOLv1 lolv1 GoPro gopro
@@
Non-generative methods
PromptIR promptir NeurIPS’23 32.02 0.952 0.013 31.98 0.924 0.115 29.56 0.888 0.087 22.89 0.847 0.296 27.21 0.817 0.250
InstructIR instructir ECCV’24 26.90 0.952 0.017 29.56 0.885 0.088 22.81 0.836 0.132 28.26 0.870 0.146
AWRaCLe awracle AAAI’25 30.81 0.979 0.013 30.56 0.904 0.088 31.26 0.908 0.068 21.04 0.818 0.146 26.78 0.820 0.248
DCPT dcpt ICLR’25 29.10 0.968 0.017 24.11 0.766 0.203 23.67 0.863 0.106 27.92 0.877 0.169
Generative methods
Diff-Plugin diffplugin CVPR’24 23.23 0.765 0.091 21.02 0.611 0.196 21.71 0.617 0.169 19.38 0.713 0.195 21.76 0.633 0.217
AutoDIR autodir ECCV’24 24.48 0.780 0.081 19.00 0.515 0.347 23.02 0.642 0.162 19.43 0.766 0.135 23.55 0.700 0.168
PixWizard pixwizard ICLR’25 21.28 0.738 0.142 21.24 0.594 0.206 21.38 0.596 0.180 15.84 0.629 0.305 20.49 0.602 0.223
RestoreVAR (Ours) 24.67 0.821 0.074 24.05 0.713 0.156 23.97 0.700 0.153 21.72 0.782 0.126 23.96 0.737 0.167
Input Diff-Plugin diffplugin AutoDir autodir PixWizard pixwizard RestoreVAR GT
@@
Table 2: Quantitative comparisons of generalization of RestoreVAR against state-of-the-art non-generative approaches on real-world images. The best result is indicated in bold.
Method Venue LHP lhprain REVIDE revide SICE sice TOLED toledpoled POLED toledpoled Average
@@
PromptIR promptir NeurIPS’23 55.813 0.334 60.277 0.390 57.651 0.512 43.320 0.249 34.604 0.259 50.333 0.349
InstructIR instructir ECCV’24 58.288 0.335 63.103 0.352 66.089 0.493 44.985 0.298 23.317 0.241 51.156 0.344
AWRaCLe awracle AAAI’25 56.767 0.361 58.295 0.351 62.067 0.512 45.932 0.310 39.493 0.264 52.511 0.360
DCPT dcpt ICLR’25 58.075 0.341 60.109 0.389 65.293 0.521 44.298 0.281 38.241 0.323 53.203 0.371
RestoreVAR 57.662 0.414 63.607 0.429 66.147 0.537 52.374 0.338 48.118 0.276 57.582 0.399
Input PromptIR promptir InstructIR instructir AWRaCLe awracle DCPT dcpt RestoreVAR GT
@@
Figure 6: Qualitative comparisons of RestoreVAR with non-generative methods on real degradations. RestoreVAR consistently achieves better results, demonstrating its generalization capability.
For testing generalization, we report MUSIQ musiq and CLIPIQA clipiqa scores in Table 2 on the real-world datasets discussed in Sec. 4.2. RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under real degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches.
@@
PromptIR promptir 2.05
InstructIR instructir 2.71
AWRaCLe awracle 2.11
DCPT dcpt 2.23
AutoDIR autodir 3.80
RestoreVAR 4.42
To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $40$ real-world scenes. We received $32$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences. More visual results are given in the supplementary.
Computational Complexity. RestoreVAR achieves substantial performance improvements over LDM-based AiOR approaches at a fraction of their computational cost. To show this, we compared RestoreVAR with Diff-Plugin, AutoDIR, and PixWizard in terms of inference steps, runtime, TeraFLOPs, and total parameter count. As shown in Table 5, RestoreVAR achieves a $\mathbf{10\times}$ speed-up over Diff-Plugin and a $\sim\mathbf{16\times}$ reduction in TFLOPs. Compared to AutoDIR and PixWizard, RestoreVAR is over $\mathbf{40\times}$ faster in inference.
These results highlight the effectiveness of RestoreVAR’s generative capabilities while being significantly faster than LDM-based methods.
Limitations and Scope for Future Work. Despite the strengths of RestoreVAR, there remains scope for improvement. First, its performance is inherently constrained by the latent refiner transformer (LRT) and the VAE decoder. While the LRT significantly improves results over using no refiner, it does not reach the upper bound set by directly decoding from ground-truth continuous latents. Exploring improved VQVAE and refiner architectures could help address this. Another promising direction is to employ our non-generative LRT in fully generative VAR models, given its strong performance for AiOR. Finally, future work can investigate how the performance of RestoreVAR scales with larger VAR backbones.
Table 4: Comparison of the computational complexity of RestoreVAR with LDM-based AiOR approaches.
Method Steps Time (s) TFLOPs Params (M)
Diff-Plugin 20 2.04 16.08 859.50
AutoDIR 100 8.477 67.80 859.50
PixWizard 60 8.247 19.27 2011.40
RestoreVAR 10 0.201 1.05 296.95
@@
4.4 Ablations
We conducted ablation studies to analyze the impact of key components of RestoreVAR.
Latent Refiner Transformer. The Latent Refiner Transformer (LRT) is critical for preserving pixel-level detail in restored images. To analyze its impact, we compare four RestoreVAR variants: (i) No refiner, (ii) HART’s diffusion refiner hart , (iii) LRT without final block outputs, and (iv) our proposed LRT. As shown in Table 5, our LRT achieves the best PSNR and SSIM, while maintaining low inference time and a low parameter count. Using no refiner yields poor PSNR/SSIM scores. Removing the last block outputs significantly reduces performance, indicating its importance as pseudo-continuous guidance for refinement. HART’s MLP diffusion-based refiner performs worse than our proposed LRT while having a significantly higher parameter count and runs $\sim\mathbf{7\times}$ slower.

### v2

@@ 3.3.1 Autoregressive Transformer Architecture
We initialize $g_{i}=0$ to retain VAR’s pretrained behavior and gradually introduce conditioning. Furthermore, we replace absolute positional embeddings in VAR with 2D Rotary Positional Embeddings (RoPE) for scaling resolution from $256\times 256$ to $512\times 512$, as RoPE is well-suited for handling varying sequence lengths (Su et al., 2024). We also remove AdaLN layers, reducing $\sim 100$M parameters with negligible impact on performance. Inference closely follows that of VAR (see Sec. 3.1), except that each scale prediction is now guided by the degraded latent. The output is a sequence of predicted indices $\{\hat{r}^{k}_{\text{pred}}\}_{k=1}^{K}$, which is then used to construct the discrete restored latent $f^{\text{pred}}_{\text{quant}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. The above steps are shown in Fig. 3(a)(ii). More architectural details are given in the supplementary.
@@ 3.3.2 Detail-Preserving Restoration
The discrete latent ($f^{\text{pred}}_{\text{quant}}$) predicted by the RestoreVAR transformer is decoded by the VQVAE to produce the restored image. However, vector-quantization and VAE decoding cause a noticeable loss of fine details in the pixel-space, leading to distorted reconstructions. This presents a major challenge for using VAR in AiOR, as the scene semantics may not be accurately preserved. To address this, we introduce VAE decoder fine-tuning on continuous latents, and a lightweight latent refinement transformer (LRT) that converts discrete latents to continuous latents for decoding.
VAE Decoder Fine-Tuning. HART (Tang et al., 2024) addressed VAE-induced distortions by fine-tuning the VAE decoder on both discrete and continuous latents. While effective for generative tasks, the VAE decoder of HART produces overly textured outputs, compromising accurate reconstruction (see supplementary). Instead, we fine-tune the decoder only on continuous latents, bypassing the quantizer. The encoder and quantizer are kept frozen, and the decoder is trained on $(f^{\text{gt}}_{\text{cont}},I_{\text{gt}})$ pairs. To avoid overly smooth outputs, we use a PatchGAN (Isola et al., 2017) discriminator (see Sec. 4.4) and optimize the decoder using pixel-wise, perceptual, and adversarial losses as
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},\vskip-4.0pt$
where $\mathcal{L}_{\text{L1}}$ is the L1 loss, $\mathcal{L}_{\text{SSIM}}$ is the SSIM loss, $\mathcal{L}_{\text{percep}}$ is the perceptual loss, $\mathcal{L}_{\text{adv}}$ is the adversarial loss and $\lambda_{i}$ are their respective weights (see Fig. 3(a)(iv)). Our fine-tuning approach yields a decoder that is well-aligned with the objectives of AiOR, achieving mean (over $1000$ samples) reconstruction PSNR/SSIM scores of 28.1428.14dB/$0.842$, outperforming both the VAR VQVAE (22.5922.59dB/$0.679$) and HART decoders (26.4826.48dB/$0.804$). Qualitative comparisons are given in the supplementary.
Refining Discrete Latents. Since the VAE decoder is fine-tuned for continuous latents, the predicted discrete latent, $f^{\text{pred}}_{\text{quant}}$, must be converted into a continuous form for decoding.
@@
While HART uses a 3737M parameter diffusion-based MLP for this, it incurs a $\sim 20\%$ inference overhead due to iterative denoising. Instead, we propose a lightweight, non-generative latent refinement transformer (LRT) that predicts a residual, which when added to $f^{\text{pred}}_{\text{quant}}$, produces a continuous latent, $\hat{f}_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$ as
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),\vskip-4.0pt$
@@ 4.1 Implementation Details
[…]
@@ 4.2 Datasets
We trained RestoreVAR for five tasks: dehazing, desnowing, deraining, low-light enhancement and deblurring. For dehazing, we used the RESIDE (Li et al., 2019) dataset comprising $72135$ training and $500$ test images. The Snow100k dataset (Liu et al., 2018) was used for desnowing, with $50000$ training and $16801$ test images (heavy subset). For deraining, we used Rain13K (Zamir et al., 2021) consisting of $13711$ training and $4298$ test images. The LOLv1 (Wei et al., 2018) dataset was used for low-light enhancement, consisting of $485$ training and $15$ test images. For deblurring, we used the GoPro (Nah et al., 2017) dataset comprising $2103$ training and $1111$ test images. We also assess generalization performance on real-world, unseen and mixed degradation datasets, namely, LHP (Guo et al., 2023) ($1000$ images), REVIDE (Zhang et al., 2021b) ($284$ images), TOLED (Zhou et al., 2021) ($30$ images), POLED (Zhou et al., 2021) ($30$ images), CDD (Guo et al., 2024) ($200$ images, mix of haze and rain), and LOLBlur (Zhou et al., 2022) ($482$ images, mix of low-light and blur). TOLED and POLED datasets contain unseen degradation of under-display camera restoration.
@@
Method Venue RESIDE Snow100k Rain13K LOLv1 GoPro
@@
[…]
PromptIR NeurIPS’23 32.02 0.952 0.013 31.98 0.924 0.115 29.56 0.888 0.087 22.89 0.847 0.296 27.21 0.817 0.250
InstructIR ECCV’24 26.90 0.952 0.017 29.56 0.885 0.088 22.81 0.836 0.132 28.26 0.870 0.146
AWRaCLe AAAI’25 30.81 0.979 0.013 30.56 0.904 0.088 31.26 0.908 0.068 21.04 0.818 0.146 26.78 0.820 0.248
DCPT ICLR’25 29.10 0.968 0.017 24.11 0.766 0.203 23.67 0.863 0.106 27.92 0.877 0.169
DFPIR CVPR’25 31.39 0.979 0.012 24.87 0.794 0.171 23.12 0.853 0.123 28.66 0.884 0.158
[…]
Input Diff-Plugin AutoDIR PixWizard RestoreVAR GT
@@
4.3 Comparisons
We compare RestoreVAR with state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), DCPT (Hu et al., 2025) and DFPIR (Tian et al., 2025). Among generative methods, we compare with the LDM-based approaches Diff-Plugin (Liu et al., 2024), AutoDIR (Jiang et al., 2023) and PixWizard (Lin et al., 2024). To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for most of our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 4.2 presents PSNR, SSIM and LPIPS scores on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR surpasses LDM-based AiOR methods at a fraction of their computational cost (inference time (s) per image)—Diff-Plugin:$2.04$s, AutoDIR: $8.477$s, PixWizard: $8.247$s and RestoreVAR: $0.201$s, highlighting the efficacy of our framework. More detailed complexity comparisons are given in the supplementary along with a derivation showing that the time complexity of VAR with maximum latent resolution $n\times n$ is $\mathcal{O}(\log n)$ lower than an LDM operating at the same latent resolution. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works (Liu et al., 2024; Jiang et al., 2023; Rajagopalan & Patel, 2024)), and assess perceptual realism through a user study.
Table 2: Quantitative comparisons of RestoreVAR against state-of-the-art non-generative approaches on real-world, unseen and mixed degradations. The best result is indicated in bold.
Method LHP REVIDE TOLED POLED LOLBlur (L ++ B) CDD (H ++ R) Average
@@ 4.3 Comparisons
PromptIR 56.780 0.366 61.191 0.459 43.218 0.281 34.536 0.303 33.693 0.166 65.895 0.483 49.219 0.343
InstructIR 58.269 0.359 63.116 0.416 44.985 0.298 23.317 0.241 40.221 0.202 65.491 0.482 49.900 0.333
AWRaCLe 57.889 0.333 59.287 0.368 44.670 0.285 40.533 0.332 38.186 0.171 66.253 0.484 51.470 0.329
DCPT 58.044 0.372 60.011 0.446 44.062 0.314 38.138 0.345 37.393 0.175 68.440 0.544 51.681 0.366
DFPIR 56.483 0.330 61.009 0.450 43.820 0.276 35.668 0.289 36.277 0.163 54.408 0.349 47.611 0.310
[…]
Input InstructIR AWRaCLe DCPT DFPIR RestoreVAR GT
@@
CDD
Figure 6: Qualitative comparisons of RestoreVAR with non-generative methods on real, unseen and mixed degradations. RestoreVAR consistently achieves better results.
For testing generalization, we report MUSIQ (Ke et al., 2021) and CLIPIQA (Wang et al., 2023) scores in Table 2 on the real-world, unseen and mixed degradation datasets discussed in Sec. 4.2.
@@
PromptIR 2.11
InstructIR 2.93
AWRaCLe 2.33
DCPT 2.42
DFPIR 2.35
AutoDIR 3.68
[…]
RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under these degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches. Due to space constraints, qualitative comparisons with PromptIR and visual results for LOLBlur are given in the supplementary. To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $50$ real-world scenes. We received $36$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences.
These results highlight the effectiveness of RestoreVAR for AiOR. We discuss limitations of RestoreVAR in the supplementary.
4.4 Ablations
@@
Latent Refiner Transformer. The Latent Refiner Transformer (LRT) is critical for preserving pixel-level detail in restored images. To analyze its impact, we compare four RestoreVAR variants: (i) No refiner, (ii) HART’s diffusion refiner, (iii) LRT without final block outputs, and (iv) our proposed LRT. As shown in Table 4, our LRT achieves the best PSNR and SSIM, while maintaining low inference time and a low parameter count. Using no refiner yields poor PSNR/SSIM scores. Removing the last block outputs significantly reduces performance, indicating its importance as pseudo-continuous guidance for refinement. HART’s MLP diffusion-based refiner performs worse than our LRT while having a much higher parameter count and runs $\sim\mathbf{7\times}$ slower.